OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (4, –2) and a radius of 9?

OpenStudy (anonymous):

You want to use the general form equation: \[(x - h)^{2} + (y - k)^{2} = r^{2}\]

OpenStudy (anonymous):

Now look at what information you were given. You have the coordinates of the center of the circle and its radius. So you will plug those into the equation. h = 4, k = -2, r = 10

OpenStudy (anonymous):

\[(x - 4)^{2} + (y + 2)^{2} = 10^2\]